# Powerformer: A Transformer with Weighted Causal Attention for Time-series Forecasting

### v1

@@ 4.3 Empirical Results
Powerformer achieves state-of-the-art performance as we compare it against other state-of-the-art models in Table 1. Powerformer achieves the best performance in 47 forecasting tasks compared to the next best model (PatchTST) which achieves 17. We note that Powerformer outperforms ETSformer in all tasks, further supporting our claim that how and where locally decaying attention weights are applied is crucial. Notably, Powerformer outperforms One-Fits-All [49] which is more than twice Powerformer’s size and is built from pre-trained GPT-2 [29].
Powerformer is sensitive to the mask type and aligns $f(t)$ with the dataset’s natural pairwise correlation distribution, which we observe as $f^{(\text{PL})}(t)$ and $f^{(\text{SPL})}(t)$ consistently outperforming $f^{(\text{BW})}_{n}(t)$. See Tables A5, A6, A7, and A8 for full $f^{(\text{PL})}(t)$, $f^{(\text{SPL})}(t)$, $f^{(\text{BW})}_{1}(t)$, and $f^{(\text{BW})}_{2}(t)$ evaluation results. The $f^{(\text{BW})}_{n}(t)$ filters resemble step functions and similarly weight all the values within the window. Alternatively, $f^{(\text{PL})}(t)$ and $f^{(\text{SPL})}(t)$ allow for longer-time couplings but significantly diminish their contribution; see Fig. 1c. The superior performance of $f^{(\text{PL})}(t)$ and $f^{(\text{SPL})}(t)$ over $f^{(\text{BW})}_{n}(t)$ further supports that the method for enforcing locality is important and that power-law correlation distributions are naturally suited to our datasets. We perform further ablation studies and report detailed results in Appendix A6.
@@
In addition to distinguishing between masks that naturally align with the data distribution, Powerformer is also sensitive to the pairwise correlation timescale. The ETTh and ETTm datasets test Powerformer’s ability to adapt the timescale $\alpha$ of $\mathbf{M}^{(\text{D})}$ for datasets sampled from the same distribution but with different temporal resolution. Since ETTh and ETTm are sampled on the hour and minute timescales, respectively, ETTm requires a slowly decaying mask to access the same information as a quickly decaying mask for ETTh. As expected, Powerformer with $f^{(\text{SPL})}(t)$ (Table A6) favors quickly decaying masks on ETTh while selecting slowly decaying masks on ETTm. This illustrates Powerformer’s sensitivity to both the temporal information timescale, the timescale covering information needed to accurately forecast, and the temporal sampling rate.
@@
To learn the dataset’s temporal information timescale, we consider a learnable $\alpha$ instead of treating it as a hyperparameter. Section A6.3 outlines method and experimental details, as well as the full results. During training, we observe that $\alpha$ consistently and monotonically drops. This is expected, as it facilitates overfitting by providing the model with extraneous information. Ultimately, we observe very little difference between learnable and constant $\alpha$ (see Table A9). This indicates that WCMHA acts as a regularizer during training, removing superfluous information to avoid overfitting and improve generalization.
4.4 Interpretability
The simplicity of Powerformer is a key strength and indicator that causal and local biases are important for time-series Transformer-based models. Previous models employ more complicated attention schemes using autocorrelations [42] or Fourier decompositions [48]; or they induce causal and local implicit biases via various mechanisms [39, 20, 19]. Some models, like FEDformer [48], purposely avoid such implicit biases. The success and simplicity of WCMHA and Powerformer over other models and their attention techniques bolsters the importance of prioritizing embeddings with causal correlation structures natural to the dataset. Moreover, Powerformer’s simplicity, along with the local and causal biases, promotes interpretability and provides insight into the dependencies of each forecast.
Figure 4: We show Powerformer’s attention score and weight distributions with MHA (dotted line) and with WCMHA (solid lines) for $f^{(\text{PL})}(t)$. The colored lines correspond to different mask decay times $(\alpha)$. These results are for the Weather dataset with a 96 prediction length and 512 input length.
Figure 5: We show the causal and local biases’ implicit and explicit effects on Powerformer’s attention score and weight distributions. The reference (dotted line) has no mask (MHA), the solid line uses WCMHA and has the mask applied, and the dashed-dotted line is the WCMHA distribution calculated before applying the mask. This result are for the Weather dataset with a 96 prediction length and 512 input length.
By leveraging Powerformer’s simplicity, we interprate it’s behavior through the attention weight and score distributions. As the power-law dampening grows (Fig. 4), an emergent bimodal distribution in attention weights appears. The score distributions (inset in Fig. 4) shift towards lower values, while the positive heavy-tail increases as the decay constant grows. This shift in scores to larger magnitude values produces an emergent bimodal distribution in the attention weights. This bimodal behavior is well aligned with our implicit biases as the right mode (larger weights) corresponds to our amplified local interactions and the left mode (smaller weights) corresponds to the diminished non-local interactions. The bimodal $\mathbf{C}^{(\text{C,D})}$ distributions in Fig. 4 indicates that $\mathbf{M}^{(\text{C,D})}$ achieves our stated goal of amplifying local couplings, removing causally inconsistent couplings, and diminishing unphysically long-time couplings.
Remarkably, we observe Powerformer further enforcing our local bias beyond $\mathbf{M}^{(\text{D})}$, and we consider this a key insight into the local bias’ importance. Because our masks are constant, Powerformer has the potential to mitigate the effects of $\mathbf{M}^{(\text{D})}$ by learning to reweight the attention scores. Instead, we observe the opposite, Powerformer further enforces our locality bias during training by imposing a bimodal attention weight distribution before $\mathbf{M}^{(\text{C,D})}$ is applied. Figure 5 shows the attention score and weight distributions with and without applying $\mathbf{M}^{(\text{C,D})}$. The dashed-dotted line in the attention weight indicates that Powerformer imposed a local bias before the locally biasing $\mathbf{M}^{(\text{C,D})}$ mask is applied. This learned effect amplifies the impact of $\mathbf{M}^{(\text{C,D})}$ as seen by the further separation in modes between the solid and dashed-dotted distributions in Fig. 5. Ultimately, Powerformer learns embeddings that further strengthen the effects of our implicit bias. We consider Powerformer’s learned amplification of our causal and local implicit biases to be strong evidence these biases are crucial in time-series forecasting.
Our local biasing power-law $\mathbf{M}^{(\text{C,D})}$ endows Powerformer with an implicit temporal multiscale structure as its embeddings change as the timescale increases. Let us consider the case when $\mathbf{M}^{(\text{D})}$ mask has a long tail (e.g., $f^{(\text{PL})}(t)$ for $\alpha\leq 0.5$ in Fig. 2). The mask weighting in this tail begins to asymptote and similarly scales the attention weights at large $\Delta t$. Summing over many similarly weighted time steps washes out high-frequency information and amplifies dynamic information on the time scale of the summation. For example, if one added a fast oscillating sinusoid to a line and summed this signal over many periods of the sinusoid, the high-frequency oscillations would cancel and one would see a linear signal. Conversely, in temporally local regions the mask quickly changes to amplify local high-frequency components. These local high-frequency components thus play a more prominent role than high-frequency components from longer windows. The mask imposes a temporal multiscale attention structure by capturing low-frequency information from longer-time summations and high frequency from local interactions. In Appendix A4, we observe this phenomena as Powerformer better predicts higher frequency contributions fast-varying signals than PatchTST.
By removing longer-term pairwise contributions, WCMHA has the potential to improve the attention complexity cost from quadratic to linear. Moreover, since multiple timescales can provide similar results (Tables A5, A6, A7, and A8), such speed improvements may come at a small cost to accuracy. Each mask type and timescale has a cutoff length $\tau$, after which the contributions from data are negligible: e.g., for $f^{(\text{PL})}(t)$ with $\alpha=1$, points beyond $\tau=100$ are diminished by two orders of magnitude and may be considered negligible. When WCMHA is only applied within $\tau$, it scales linearly O$(\tau T)$ with input sequence length $T$, instead of quadratically (O$(T^{2})$) like MHA. Because $\alpha$ varies between datasets and can be lowered to trade speed for small changes in accuracy, we expect a large variance in potential speedup times, and so we do not provide results here. For example, $f^{(\text{PL})}(t)$ with $\alpha=1$ provides a factor of 3 and 5 speedup for typical look-back window of 336 and 512, respectively.
WCMHA leverages the strengths of both statistical models and MHA. The induced power-law decay in attention weights, and thus temporal correlations, resembles previous statistical models. However, the WCMHA attention mechanism is not fully overwhelmed by this decay and still amplifies important longer-time correlations, as shown in Fig. 1b and Fig. 1c where the attention mask clearly selects regions of strong correlations near and past the tail of the power-law decay. We believe this balance between power-law temporal correlations and maintaining WCMHA’s ability to attend to highly correlated input is key to WCMHA’s success.
@@ 5 Conclusion
We develop Powerformer to impose and investigate the importance of causal and local biases for Transformer-based time-series models. Powerformer outperforms state-of-the-art time series models on standard tasks often with a simpler architecture. Powerformer’s simplicity promotes interpretability studies that elucidate the effects of our biases. Powerformer strengthens our local bias by learning to decrease attention scores for long-time interactions.
In addition to outperforming previous models on forecasting accuracy, Powerformer also has the potential to significantly increase attention efficiency. The decaying nature of $f(t)$ determines a cutoff time after which $\mathbf{M}^{(\text{D})}$ effectively removes contributions. Evaluating within this cutoff time will result in O($\tau T$) rather than O($T^{2}$) complexity.
@@ Acknowledgements
NBE would like to acknowledge the Laboratory Directed Research and Development Program of Lawrence Berkeley National Laboratory under U.S. Department of Energy Contract No. DE-AC02-05CH11231, and NERSC (DE-AC02-05CH11231) for providing compute resources.
Appendix
Appendix A1 Butterworth Filter
@@
The Butterworth filter [6] is often used in signal processing for low-, high-, and band-pass filters. It is designed to be optimally flat within the passband and sometimes referred to as the “maximally flat filter.” The motivation behind it’s derivation is to have equal sensitivity to frequency modes within the passband. We leveraged the Butterworth filter to equally weight temporal measurements within the lookback window. The analog Butterworth filter gain is given by
$f^{\text{BF}}_{n}(z)=\frac{-1}{\sqrt{1+\left(\frac{z}{t_{\text{c}}}\right)^{2n}}},$ (A1)
@@
where $t_{c}$ is the critical time that sets the width of the filter and the order $n$ determines how fast the gain decays after $t_{c}$. In this work we use the digital Butterworth filter gain. Figure A1 shows the $f^{(\text{BW})}_{1}(t)$ and $f^{(\text{BW})}_{2}(t)$ contributions on the attention score and weights for varying decay constants.
@@
Figure A1: We show the effects of the Butterworth filter masks $\left(f^{(\text{BW})}_{n}(t)\right)$ for orders 1 (dashed lines) and 2 (solid lines), while varying the critical time. Panel (a) shows the effects on the attention scores and Panel (b) shows the corresponding effects on the attention weights after applying Softmax to $f^{(\text{BW})}_{n}(t)$.
A1.1 Digital Filter Gain
@@
A1.2 Code to Calculate the Gain
@@
Appendix A2 Datasets
We evaluate Powerformer on 7 real-world datasets that have become standard public benchmarks for time-series forecast tasks. The datasets can be found and downloaded in Ref. [42], or individually at references below.
Table A1: We provide the number of variables and the number of timesteps for each dataset.
@@
ETT [^1] https://github.com/zhouhaoyi/ETDataset [47] provides 7 measurements of the electrical transformer power load (including oil temperature) between July 2016 and July 2018. This is done over 2 regions in China at 2 different sampling rates (hourly and every 15 minutes), resulting in 4 separate datasets (ETTh1, ETTh2, ETTm1, ETTm2). We show the pairwise correlation dependence in Fig. A5.
@@
Weather [^2] https://www.bg-ena.mpg.de/wetter/ [42] provides 21 meteorological measurements (air temperature, air pressure, humidity, precipitation, etc.) collected in Germany over the whole of 2020. These measurements are recorded every 10 minutes. We show the pairwise correlation dependence in Fig. A4.
@@
Electricity [^3] https://archive.ics.uci.edu/ dataset/321/electricityloaddiagrams20112014 [42] provides the electricity consumption (kWh) of 321 consumers from 2012 to 2014. These measurements are sampled every hour. We show the pairwise correlation dependence in Fig. A4.
@@
Traffic [^4] http://pems.dot.ca.gov [42] provides occupancy rates on San Francisco Bay Area freeways from 826 sensors. This data comes from the California Department of Transportation and is sampled hourly. We show the pairwise correlation dependence in Fig. A2.
@@
Figure A2: We show the Weather dataset’s pairwise correlations as a pairwise separation function. Each colored line represents a different variable in the dataset.
@@
Figure A3: We show the Electricity dataset’s pairwise correlations as a pairwise separation function. Each colored line represents a different variable in the dataset.
@@
Figure A4: We show the Weather dataset’s pairwise correlations as a pairwise separation function. Each colored line represents a different variable in the dataset.
@@
Figure A5: We show the ETT datasets’ pairwise correlations as a pairwise separation function. Each colored line represents a different variable in the dataset.
Appendix A3 Architecture Details
A3.1 Transformer
@@
We use a vanilla encoder-decoder Transformer architecture for all Transformer experiments. The decoder input is the last 48 time-steps in the sequence concatenated by an array of zeros with a length of the prediction window. For our baseline experiment, the decoder self-attention has a causal mask, which is common in Transformer architectures, while the encoder self-attention and decoder cross-attention do not have masks. Our WCMHA variant of the Transformer applies WCMHA to both the encoder and decoder self-attention, but not the decoder cross-attention. We do not apply WCMHA to the decoder cross-attention because this process is already causal as it forecasts future points based on the given sequence. We present the Transformer hyperparameters used in this work in Table A2.
A3.2 PatchTST
@@
Here, we review PatchTST’s patching mechanism, as PatchTST is otherwise similar to a traditional Transformer with a linear readout head that simultaneously predicts all future points. For a detailed description of PatchTST, see Ref. [27]. Let us consider an input time-series $\mathbf{X}$ with dimensions $[B,T_{\text{seq}},D]$, where $B$ is batch size, $T_{\text{seq}}$ is the time interval, and $D$ is the number of input dimensions. The multivariate input is first normalized to zero mean and unit variance along the time axis and flattened into $D$ univariate inputs, resulting in a dimensionality of $[B\times D,T_{\text{seq}}]$. PatchTST batches these time series via a convolution with $N$ filters, a stride $s$, and patch size $p$, resulting in a dimensionality of $[B\times D,P,N]$. Here, $P=(T_{\text{seq}}-p)//s+1$ is the number of patches, and ///// / is integer division. At this point, the patched data is fed into the transformer where a positional embedding is added and the patched embeddings are further contextualized by vanilla MHA [36]. The linear output head takes the $[B\times D,P,N]$ output of the Transformer, flattens the last two dimensions, and applies a linear matrix multiplication (of shape $[P\times N,T_{\text{pred}}]$) to predict the output time series of length $T_{\text{pred}}$. The linear head returns a matrix of shape $[B\times D,T_{\text{pred}}]$. This matrix is reshaped into the original multivariate dimensionality $([B,T_{\text{pred}},D])$ and renormalized. We note that the patching mechanism is performed before applying the Transformer and is independent of it.
A3.3 Powerformer

### v2

@@ Results
Compared to our benchmarks (Table 1), Powerformer achieves the best performance against more complicated (FEDformer (Zhou et al., 2022), ETSformer (Woo et al., 2022), One-Fits-All (Zhou et al., 2023)) models. Powerformer achieves the best performance in 46 forecasting tasks compared to the next best model (PatchTST) which achieves 9. We note that Powerformer outperforms ETSformer, which employs an exponential temporal decay without attention. This result reinforces the importance of correctly balancing RBCA’s recency bias with attentions ability to learn data-dependent correlations. Notably, Powerformer outperforms One-Fits-All (Zhou et al., 2023), which is a pre-trained GPT-2 model more than twice Powerformer’s size.
Qualitatively, Powerformer outperforms the next best model (PatchTST) by capturing high-frequency variability. Powerformer better models sharper changes in oscillations and deviations from a generic sinusoidal signal, as demonstrated in Appendix S4. This behavior can be described in the context of the power-law’s heavy tail, which induces a multi-scale attention structure. During attention, the embeddings in the long tail will be summed together with relatively similar weights since the heavy tail and the unbiased attention weights (Fig. 2) are relatively constant. This summation removes short-lived input sequence frequency components with periods less than the tail’s length. This is because the summation covers multiple, if not many, periods of these frequency components, causing them to cancel out. However, long-term signals with periods greater than the tail length will be captured. Therefore, the tail highlights longer-term dynamics in the input sequence. Conversely, the quickly varying power-law near 0 amplifies a shorter context window. This shorter window highlights local higher-frequency components by reducing the period at which the signal is washed out in the summation. Again, striking a balance between attending to long-term sinusoidal signals and quickly decaying exponential pair-wise distributions. Therefore, the power-law recency bias endows Powerformer with an implicit temporal multiscale structure: the tail focuses on low-frequency signal while high-frequency components come from the most recent signal.
Powerformer underperforms for the Traffic dataset and outperforms by a smaller margin in the Electricity dataset due to these datasets’ highly periodic nature. As highlighted, Powerformer can capture short and long-term sinusoidal signals, but long and constant attention windows – used in all the other benchmarks – are better suited to highly-periodic signals. In cases where the data is highly periodic, like Traffic (Fig. S5) and Electricity (Figs. S7 and S9), these systems can often be effectively modeled by simpler Fourier or statistical methods without requiring the much larger training and evaluation overhead of Transformers. Powerformer is built as a general-purpose time-series method to handle challenging non-stationary dynamics.
Powerformer is sensitive to the recency bias and aligns $f(t)$ with the dataset’s specific pairwise correlation. To test this, we evaluated the first 5 datasets on $f^{(\text{BW})}_{n}(t)$, which induces a flat pairwise correlation structure. The $f^{(\text{PL})}(t)$ and $f^{(\text{SPL})}(t)$ biases consistently outperform $f^{(\text{BW})}_{n}(t)$, see Tables S5, S6, S7, and S8 for the full evaluation results of $f^{(\text{PL})}(t)$, $f^{(\text{SPL})}(t)$, $f^{(\text{BW})}_{1}(t)$, and $f^{(\text{BW})}_{2}(t)$, respectively. The superior performance of $f^{(\text{PL})}(t)$ and $f^{(\text{SPL})}(t)$ over $f^{(\text{BW})}_{n}(t)$ demonstrates Powerformer’s sensitivity to the dataset’s correlation structure.
@@
In addition to distinguishing between biases that naturally align with the data distribution, Powerformer is also sensitive to the dataset’s correlation time scale. The ETTh and ETTm datasets test Powerformer’s ability to adapt the timescale $\alpha$ of $\mathbf{M}^{(\text{R})}$ for datasets sampled from the same distribution but with different temporal resolution. Since ETTh and ETTm are sampled on the hour and minute timescales, respectively, ETTm requires a slowly decaying mask to access the same information as a quickly decaying mask for ETTh. As expected, Powerformer with $f^{(\text{SPL})}(t)$ (Table S6) favors fast decaying masks on ETTh and slowly decaying masks on ETTm. This illustrates Powerformer’s sensitivity to adjust its timescale to cover the same information content in each dataset.
4.4 Ablation
We ablate Powerformer to determine the effects of the causal mask ($\mathbf{M}^{(\text{C})}$) and the addition of the recency bias ($\mathbf{M}^{(\text{R})}$). Table S8 compares the performance of Powerformer with MHA (without $\mathbf{M}^{(\text{C,R})}$), with only the causal mask ($\mathbf{M}^{(\text{C})}$), and with RBCA ($\mathbf{M}^{(\text{C,R})}$). We first observe that causal-only attention ($\mathbf{M}^{(\text{C})}$) outperforms MHA, validating the importance of our causal inductive bias. In all but one case, RBCA ties or outperforms causal-only attention.
@@ 4.4 Ablation
To learn the temporal information timescale, we make $\alpha$ learnable instead of treating it as a hyperparameter. Section S8.1 outlines the experimental details and full results. During training, $\alpha$ drops consistently and monotonically. This is expected as a small magnitude $\alpha$ facilitates overfitting by providing the model with extraneous (long-term) information. To keep $\alpha$ from converging to 0, we initialize $\alpha$ with the hyperparameter tuning values and quickly decay the learning rate. The learned $\alpha$ values decrease in magnitude, but we observe very little difference between the learnable and constant $\alpha$ (Table S10). This indicates that RBCA regularizes the model, removing superfluous information to avoid overfitting and improve generalization.
@@ 5 CONCLUSION
We develop RBCA to test the importance of causal and recency biases in Transformers, as they are applied to causal and locally correlated NLP and time-series tasks. Some recency biases ($f^{(\text{PL})}(t)$) are sufficient requirements to imbue Transformer with advantageous RNN capabilities: read and write to memory, and ignore input information. However, common NLP recency biases ($f^{(\text{E})}(t)$ and $f^{(\text{BW})}_{n}(t)$) diminish these capabilities. This indicates that locality biases are important for Transformers to gain important capabilities, but the bias’ implementation matters.
We develop Powerformer to impose and investigate the importance of causal and recency biases for Transformer-based time-series models. Powerformer, with its simple architecture, outperforms larger and more complex models. Notably, Powerformer strengthens the causal and recency biases by enforcing a power-law decay before the $\mathbf{M}^{(\text{C,R})}$ is applied. This is strong evidence that the causal and recency biases are important for time-series problems, as Powerformer is essentially removing information at training time that could be used to overfit.
@@ Acknowledgements
The authors thank Dongwei Lyu for her help in running the flip flop experiments to test the RBCA Transformer glitching. NBE would like to acknowledge the Laboratory Directed Research and Development Program of Lawrence Berkeley National Laboratory under U.S. Department of Energy Contract No. DE-AC02-05CH11231, and the Scientific Discovery through Advanced Computing (SciDAC) program, under Contract Number DE-AC02-05CH11231 at Berkeley Lab, and NERSC (DE-AC02-05CH11231) for providing compute resources.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes]
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes]
    2. (b)
      The license information of the assets, if applicable. [Yes]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Yes]
    4. (d)
      Information about consent from data providers/curators. [Yes]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
Supplementary Materials
S1 LOCALITY INDUCING DECAY FUNCTIONS
S1.1 Power-Law Functions
The two power-law-based masks $f^{(\text{PL})}(t)$ and $f^{(\text{SPL})}(t)$ apply a power-law decay to the attention weights and scores, respectively. Here, we provide intuition for each mask type a well as illustrate the masks used in our evaluation.
We consider inducing a multiplicative power-law decay to the attention weights ($f^{(\text{PL})}(t)$) to be the most natural way of inducing locality:
$f^{(\text{PL})}(t)=-\alpha\,\log(t).$ (S1)
The attention weights are analogous to correlation coefficients. Therefore, since power-law distributions show up in correlation structures, we would like the attention weights to follow a similar power-law distribution. When training Transformer and Powerformer with $f^{(\text{PL})}(t)$ we used decay times of $\alpha\in[0.1,0.25,0.5,0.75,1.0]$. Figure S2 shows the additive attention mask and the resulting multiplicative effects on the attention weights.
Figure S1: The power-law masks ($f^{(\text{PL})}(t)$) used when evaluating RBCA on Transformer and when evaluating Powerformer. Panel (a) shows the additive mask, (b) shows the multiplicative effects on the attention weights.
Figure S2: The similarity power-law masks ($f^{(\text{SPL})}(t)$) used when evaluating RBCA on Transformer and when evaluating Powerformer. Panel (a) shows the additive mask, (b) shows the multiplicative effects on the attention weights.
We additionally apply the power-law directly to the similarity score:
$f^{(\text{SPL})}(t)=-(t)^{-\alpha}.$ (S2)
During the attention calculation, we get an exponential of this power-law, which results in steep initial decay and a long tail. When training Transformer and Powerformer with $f^{(\text{SPL})}(t)$ we used decay times of $\alpha\in[0.1,0.5,1.0,2.0]$. Figure S2 shows the additive attention mask and the resulting multiplicative effects on the attention weights.
Combining the $f^{(\text{PL})}(t)$ and $f^{(\text{SPL})}(t)$ masks, we cover decays starting from a relatively flat mask ($f^{(\text{PL})}(t)$ with $\alpha=0.5$) to a sharply decaying ($f^{(\text{SPL})}(t)$ with $\alpha=2.0$). To illustrate our coverage we plot $f^{(\text{PL})}(t)$ and $f^{(\text{SPL})}(t)$ together in Fig. S3.
Figure S3: The $f^{(\text{PL})}(t)$ and $f^{(\text{SPL})}(t)$ mask (a) and their effects on the weights (b) are shown to illustrate the coverage of decays used when assessing Powerformer.
S1.2 Butterworth Filter
@@
The Butterworth filter (Butterworth, 1930) is often used in signal processing for low-, high-, and band-pass filters. It is designed to be optimally flat within the passband and sometimes referred to as the “maximally flat filter.” The motivation behind it’s derivation is to have equal sensitivity to frequency modes within the passband. We leveraged the Butterworth filter to equally weight temporal measurements within the lookback window. The analog Butterworth filter gain is given by
$f^{(\text{BW})}_{n}(z)=\frac{-1}{\sqrt{1+\left(\frac{z}{t_{\text{c}}}\right)^{2n}}},$ (S3)
@@
where $t_{c}$ is the critical time that sets the width of the filter and the order $n$ determines how fast the gain decays after $t_{c}$. In this work we use the digital Butterworth filter gain. Figure S4 shows the $f^{(\text{BW})}_{1}(t)$ and $f^{(\text{BW})}_{2}(t)$ contributions on the attention score and weights for varying decay constants.
@@
Figure S4: We show the effects of the Butterworth filter masks $\left(f^{(\text{BW})}_{n}(t)\right)$ for orders 1 (dashed lines) and 2 (solid lines), while varying the critical time. Panel (a) shows the effects on the attention scores and Panel (b) shows the corresponding effects on the attention weights after applying Softmax to $f^{(\text{BW})}_{n}(t)$.
S1.2.1 Digital Filter Gain
@@
S1.2.2 Code to Calculate the Gain
@@
S2 DATASETS
We evaluate Powerformer on 7 real-world datasets that have become standard public benchmarks for time-series forecast tasks. The datasets can be found and downloaded in Ref. (Wu et al., 2021), or individually at references below.
Table S1: We provide the number of variables and the number of timesteps for each dataset.
@@
ETT [^1] https://github.com/zhouhaoyi/ETDataset (Zhou et al., 2021) provides 7 measurements of the electrical transformer power load (including oil temperature) between July 2016 and July 2018. This is done over 2 regions in China at 2 different sampling rates (hourly and every 15 minutes), resulting in 4 separate datasets (ETTh1, ETTh2, ETTm1, ETTm2). We show the pairwise correlation dependence in Fig. S8.
@@
Weather [^2] https://www.bg-ena.mpg.de/wetter/ (Wu et al., 2021) provides 21 meteorological measurements (air temperature, air pressure, humidity, precipitation, etc.) collected in Germany over the whole of 2020. These measurements are recorded every 10 minutes. We show the pairwise correlation dependence in Fig. S7.
@@
Electricity [^3] https://archive.ics.uci.edu/ dataset/321/electricityloaddiagrams20112014 (Wu et al., 2021) provides the electricity consumption (kWh) of 321 consumers from 2012 to 2014. These measurements are sampled every hour. We show the pairwise correlation dependence in Fig. S7.
@@
Traffic [^4] http://pems.dot.ca.gov (Wu et al., 2021) provides occupancy rates on San Francisco Bay Area freeways from 826 sensors. This data comes from the California Department of Transportation and is sampled hourly. We show the pairwise correlation dependence in Fig. S5.
@@
Figure S5: We show the Weather dataset’s pairwise correlations as a pairwise separation function. Each colored line represents a different variable in the dataset.
@@
Figure S6: We show the Electricity dataset’s pairwise correlations as a pairwise separation function. Each colored line represents a different variable in the dataset.
@@
Figure S7: We show the Weather dataset’s pairwise correlations as a pairwise separation function. Each colored line represents a different variable in the dataset.
@@
Figure S8: We show the ETT datasets’ pairwise correlations as a pairwise separation function. Each colored line represents a different variable in the dataset.
S3 ARCHITECTURE DETAILS
S3.1 Transformer
@@
We use a vanilla encoder-decoder Transformer architecture for all Transformer experiments. The decoder input is the last 48 time-steps in the sequence concatenated by an array of zeros with a length of the prediction window. For our baseline experiment, the decoder self-attention has a causal mask, which is common in Transformer architectures, while the encoder self-attention and decoder cross-attention do not have masks. Our RBCA variant of the Transformer applies RBCA to both the encoder and decoder self-attention, but not the decoder cross-attention. We do not apply RBCA to the decoder cross-attention because this process is already causal as it forecasts future points based on the given sequence. We present the Transformer hyperparameters used in this work in Table S2.
S3.2 PatchTST
@@
Here, we review PatchTST’s patching mechanism, as PatchTST is otherwise similar to a traditional Transformer with a linear readout head that simultaneously predicts all future points. For a detailed description of PatchTST, see Ref. (Nie et al., 2023). Let us consider an input time-series $\mathbf{X}$ with dimensions $B,T_{\text{seq}},D]$, where $B$ is batch size, $T_{\text{seq}}$ is the time interval, and $D$ is the number of input dimensions. The multivariate input is first normalized to zero mean and unit variance along the time axis and flattened into $D$ univariate inputs, resulting in a dimensionality of $B\times D,T_{\text{seq}}]$. PatchTST batches these time series via a convolution with $N$ filters, a stride $s$, and patch size $p$, resulting in a dimensionality of $B\times D,P,N]$. Here, $P=(T_{\text{seq}}-p)//s+1$ is the number of patches, and //// is integer division. At this point, the patched data is fed into the Transformer where a positional embedding is added and the patched embeddings are further contextualized by vanilla MHA (Vaswani et al., 2017). The linear output head takes the $B\times D,P,N]$ output of the Transformer, flattens the last two dimensions, and applies a linear matrix multiplication (of shape $P\times N,T_{\text{pred}}]$) to predict the output time series of length $T_{\text{pred}}$. The linear head returns a matrix of shape $B\times D,T_{\text{pred}}]$. This matrix is reshaped into the original multivariate dimensionality $([B,T_{\text{pred}},D])$ and renormalized. We note that the patching mechanism is performed before applying the Transformer and is independent of it.
S3.3 Powerformer